# references.bib

### v1

@@ 2.2 Angle-based generalised coordinates
    As the first step, we place the bottom left corner elementary cell of the lattice, which has six degrees of freedom (Fig. 3.B, box 1). The six parameters we choose to describe this elementary cell are the location of the lattice and the orientation of the bottom and top beams and levers.
@@
    Then we add the first column, i.e. we stack up elementary cells on top of the first one (Fig. 3.B, box 2). Each elementary cell in this column has three degrees of freedom. Here, we choose the orientation of bottom and top lever as well as the orientation of the upper beam as parameters.
@@
    We then add the first row, i.e. we string together elementary cells to the right of the first one (Fig. 3.B, box 3). These elementary cells only have two degrees of freedom, which we choose to be the orientation of the bottom beam and lever. In this case, the midpoint of the upper beam ($\mathrm{P^{\prime}}$) is given by the (analytic) solution of a quadratic equation representing the Totimorphic condition (Fig. 3.C), see Section 4.2.2 for details.
@@
    Then we fill in the remaining elementary cells row-wise from left to right (Fig. 3.B, box 4). All these elementary cells have only one degree of freedom, which we choose to be represented by the orientation of the lower lever. The midpoint of the upper beam is again given by the solution of a quadratic equation.
@@
For a detailed mathematical description, see Section 4.2. Although we parametrise the lattice using beam and lever angles here, many other parametrisations are possible. From the quadratic equations, we further derive upper and lower bounds for the lever angles, allowing levers to be pushed or dragged by other lattice elements (Section 4.2.3). For lattices in two dimensions, additional conditions are derived to avoid configurations with overlapping lattice elements (Section 4.2.3). Models of lattices in three dimensions are constructed analogously (Section 4.3).
With this approach, a Totimorphic lattice is represented by a set of $df$ parameters $\mathcal{P}$, $|\ \mathcal{P}\ |=df$, which in our case is the location of the lattice’s bottom left corner in space as well as the angles of levers and beams in the lattice. From this set of parameters, there exists a differentiable function $f_{\mathrm{T}}$ which maps to the points of each elementary cell in the lattice (lever and beam endings), i.e. $\mathcal{R}=f_{\mathrm{T}}(\mathcal{P})$, where $\mathcal{R}$ is a set of vectors containing the coordinates of each point in the lattice – we call this the Totimorphic model in the following. Invalid lattice configurations are given by values of the generalised coordinates $\mathcal{P}$ for which no solution of $f_{\mathrm{T}}$ exists, which can be checked using simple criteria. Note that, similar to how in a neural network outputs are calculated layer by layer, points are calculated elementary cell by elementary cell; with points of later elementary cells (i.e. higher row or further to the right) depending on the configurations of elementary cells that came before. This way, a point in, e.g., the uppermost layer can be moved into a desired target position by changing the angles of elementary cells in the bottom layer.
To change the configuration of the lattice to one with some desired property (e.g. a desired stiffness), we minimise an objective-dependent cost function that measures how well the current configuration satisfies the desired property. In most cases, we express the cost function using the point coordinates obtained from the Totimorphic model. What is actually being optimised, however, are the parameters $\mathcal{P}$ used to characterise the lattice; generally these parameters number far less than the physical lattice coordinates. By tracking intermediate configurations, we also get the trajectory through parameter space connecting the initial and the final lattice configuration – which is a major benefit compared to previously employed methods [chaudhary2021totimorphic]. In this work, we exclusively use automatic differentiation (i.e. gradient descent) as our optimisation algorithm of choice to guarantee scalability, but other techniques can be used as well. To demonstrate the introduced framework, we focus on use cases where effective properties of the lattice have to be optimised, i.e. global properties of the lattice structure [^3] Typically, in these cases no specific target locations exist for individual lattice points.. As a first proof of concept, we explore reconfigurable lattice structures with adjustable mechanical properties, namely the Poisson’s ratio which characterises how a mechanical lattice responds when compressed [^4] As shown in [dold2023differentiable], the exact same approach can be used to optimise stiffness as well, which is arguably simpler than optimising the Poisson’s ratio. For this reason, we only focus on the Poisson’s ratio in this work.. As a second proof of concept, we demonstrate a deformable telescope mirror made from a Totimorphic lattice capable of continuously adjusting the location of its focal point.
@@
2.3 Continuous inverse design of lattice structures
In the following, we use the introduced framework to continuously change the Poisson’s ratio of a mechanical lattice (on the length scale of cm) through continuous changes of the Totimorphic configuration. The used approach is based on the work and simulation code presented in [dold2023differentiable], where effective mechanical properties of disordered lattices were optimised by adjusting the lattice topology using a differentiable framework.
@@ 2.3 Continuous inverse design of lattice structures
To predict the mechanical properties of a Totimorphic lattice, we use the direct stiffness method (see Section 4.4) to set up a compression experiment (see Section 4.4.2). In a compression experiment, the lattice is glued between surfaces and slowly compressed. During compression, the response of the lattice to this load is measured, for example, the displacement of nodes in the lattice – from which mechanical properties such as effective stiffness and Poisson’s ratio are derived. Here, we restrict the work to the effective Poisson’s ratio, which measures how a material expands on average horizontally when compressed vertically (see Section 4.4.3). A positive Poisson’s ratio (like honeycomb tiles) means the material widens, while a negative one means it narrows (also known as auxetic). A Poisson’s ratio of zero consequently means that the material does not change its width during compression.
Figure 4: Inverse design of mechanical properties of Totimorphic lattices. (A) Schematic of the inverse design pipeline. (B) During optimisation, we obtain a continuous trajectory through configuration space (here represented using the first principal component of the angles representing the lattice) connecting the initial configuration ($\nu=0$) and the two final configurations ($\nu=\pm 0.5$). Since the lattice always fulfills the Totimorphic condition, it passes through all intermediate Poisson’s ratios while morphing. (C) Illustration of the final design for $\nu=0.5$ (left) and $\nu=-0.5$ (right), shown without (left) and with load (right, red arrows).
@@
For the direct stiffness approach, we only consider beam and lever elements in the lattice, as they are the major supporting elements of the lattice. We model these using Generalized Euler-Bernoulli beam elements [ochsner2018finite] (see Section 4.4.1). Furthermore, we assume that joints in the Totimorphic lattice are locked during the compression experiment, meaning that the parameters $\mathcal{P}$ are not changed while applying a load. To do a single step of the compression experiment, we first calculate the stiffness matrix from the lattice points $\mathcal{R}$. Using the stiffness matrix, we can then derive the displacement of each node in the lattice resulting from the compression by solving a linear equation. We repeat this process several times and note down the horizontal and vertical strain (mean length change). Finally, the Poisson’s ratio, which is defined as the slope of the strain-strain curve, is obtained via linear regression. This process can be formalized as a single differentiable function $f_{\mathrm{DS}}$ which takes as input the coordinates of the lattice nodes and returns the Poisson’s ratio $\nu=f_{\mathrm{DS}}\left(f_{\mathrm{T}}(\mathcal{P})\right)$. The trajectory in parameter space (i.e., parameter changes $\Delta\mathcal{P}$) to reconfigure the lattice until it has a desired target Poisson’s ratio $\nu^{\mathrm{tgt}}$ is then obtained by descending the gradient of the cost function $C$,
@@
$\Delta$ $\mathcal{P}\propto-\nabla_{\mathcal{P}}C(\mathcal{P})\,,$ (2)
@@
where $\|\cdot\|_{1}$ is the L1 norm. Further details about the mathematical formulation of the direct stiffness method and the Poisson’s ratio can be found in Section 4.4. The whole optimisation pipeline is illustrated in Fig. 4.A.
@@
We demonstrate this approach for a $4\times 4$ Totimorphic lattice with beam length $l_{\text{b}}=1\,\mathrm{cm}$, beam cross-area $A=4\,\mathrm{mm}^{2}$, Young’s modulus of the beam material of $E=1\,\mathrm{GPa}$ (around the order of magnitude of polymers), and second moment of area $I=\frac{A^{2}}{12}$ (representing a square-shaped cross-sectional area). The initial lattice has a Poisson’s ratio of approximately zero, which is not surprising as it resembles a square lattice. From this initial configuration, we perform gradient descent to get to a configuration with Poisson’s ratio $\nu^{\mathrm{tgt}}=\pm 0.5$. This results in two trajectories starting from the initial configuration and ending at a configuration with the desired Poisson’s ratio, shown in Fig. 4.B using the first principal component of the lattice parameters $\mathcal{P}$. The final lattice configurations and their response to applying load is shown in Fig. 4.C. In case of the configuration with positive Poisson’s ratio (Fig. 4.C, left), large pockets were formed that widen during compression. In case of negative Poisson’s ratio (Fig. 4.C, right), the resulting lattice geometry is composed of square-shaped elements (one beam and two levers surrounded by springs) that pivot inwards when compressed from above. This effect is most pronounced in the lower half of the lattice structure. Similar designs were obtained for larger lattices, e.g., $8\times 8$ trained to achieve even higher Poisson’s ratios (see supplemental information). It is worthy noticing that every intermediate configuration obtained during optimisation is a legitimate Totimorphic lattice, and hence we are capable of continuously adjusting the Poisson’s ratio by continuously changing the lattice angles. For an animation of the reconfiguration process and the compression experiment (after reconfiguration), see Video S1a,b and Video S2a,b.
@@
2.4 Deployable and reconfigurable large-scale structures
@@
In the following, we envision that the Totimorphic lattice (on the length scale of meters) acts as a skeleton for, e.g., a graphene sheet that forms the reflective surface of a primary mirror of a space telescope [rabien2023adaptive]. We assume a $6\times 6$ Totimorphic lattice with $l_{\text{b}}=1$m in its flat surface configuration. By only changing the lever angle of all elements in the lattice (i.e., setting them all to the same value), we can switch from the flat surface configuration to a collapsed one (and vice versa), greatly reducing the surface area and volume required to launch the Totimorphic structure (see Fig. 5). The reduction in surface area is given by $\sin\theta_{\mathrm{min}}$, where $\theta_{\mathrm{min}}$ is the minimum angle the levers can be set to physically. Consequently, for launching, the lattice is configured in its collapsed configuration and for deployment, the lever angles are slowly increased to unfold the lattice into a flat surface.
@@
After deployment, the lattice configuration is continuously adjustable to change the optical properties of its mirror surface. In our case, the goal is to focus the light onto a secondary optical element such as a detector or a secondary mirror located at point $\boldsymbol{f}^{\mathrm{tgt}}$, which we call the target focal point in the following (Fig. 6.A). We utilise a similar approach as for the material inverse design use-case, meaning that we construct a differentiable framework that allows us to find suitable lattice configurations (and trajectories connecting them) using gradient-based optimisation. First, for simplicity, we assume that each unit cell reflects light at the mid-point of the lever in simulations (i.e., the mid-point between $\mathrm{C}$ and $\mathrm{P}$ as well $\mathrm{C}$ and $\mathrm{P^{\prime}}$, see Fig. 6.A). In addition, we assume that the object to be imaged is far away such that the incoming light is parallel. Given an incidence direction $\boldsymbol{I}$ ($\|\boldsymbol{I}\|_{2}=1$, where $\|\cdot\|_{2}$ is the Euclidean norm) of the incoming light, the direction of the reflected light originating from each unit cell $(ij)$ is calculated using ray-based optics (see Section 4.5.1). Subsequently, we obtain the point where each reflected light ray is closest to the target focal point (see Section 4.5.2). The whole calculation can again be formulated as a series of differentiable operations, which we summarise in two functions: the Totimorphic model ($f_{\mathrm{T}}$), and the optics part $f_{\mathrm{O}}$, such that $\boldsymbol{F}=f_{\mathrm{O}}(f_{\mathrm{T}}(\mathcal{P}))$ is a list of vectors containing the closest point to $\boldsymbol{f}^{\mathrm{tgt}}$ on each reflected light ray.
Figure 6: Deployment and reconfiguration of a Totimorphic telescope mirror. (A) Schematic of the studied setup. Given an object to be imaged, the unit cells of the Totimorphic lattice are used as mirror elements to focus the light on an imaging plane. By changing the lattice configuration, the focal point and focal length are adjusted.(B) Reconfiguration of the lattice to several different focal lengths, starting from a flat configuration. A few parameters as well as the objective-dependent cost are shown. (C) Lattice configuration and images received when using the lattice as a mirror, for configurations featuring different focal lengths.
@@
$\Delta\mathcal{P}\propto$ $-\nabla_{\mathcal{P}}C(\mathcal{P})\,,$ (4)
[…]
@@
As a proof of concept, we use the Totimorphic lattice to image a real object; specifically, a photograph of Pluto at the same distance and angular extension of the Moon (as seen from Earth), see Fig. 6.A. To construct the image, we record where light coming from each pixel of the image hits the imaging plane after being reflected by the unit cells of the lattice – imitating the real process of image generation in a telescope (see Section 4.5.4 for details). The resulting image closely resembles the original object, resolving small details such as contours of icy regions on Pluto (Fig. 6.C). Moreover, as expected, decreasing the focal length of the Totimorphic mirror increases its field of view; enabling adaptive zoom-in and out by reconfiguring the lattice.
Although we only demonstrate reconfiguration for a few discrete focal lengths, in principle, the telescope can achieve a continuum of focal points by successively adjusting the target focal length. In addition, the telescope is not limited to focal points that are on the central axis of the telescope surface, as shown in Fig. 7.A-B where we guide the focal point on a circle around this axis using the above gradient-based optimisation.
The presented approach also allows to use reconfiguration of the lattice surface to compensate for defects introduced by mirror damage, e.g., through micrometeoroids [rigby2023science]. We illustrate this by adding a single defect to one of the lattice’s unit cells (selected randomly), which is modelled by adding a deflection (randomly determined but static) to the light rays reflected from this cell (Fig. 7.C; see Section 4.5.3 for details). As before, we perform gradient descent on the respective cost function to find a lattice configuration that compensates for this defect (Fig. 7.C-D). This is studied for increasing degrees of damage, i.e., for increasing amounts of deflection. For all cases, the reconfiguration improves the performance of the mirror – although for more extreme damage, we are not able to return to the baseline performance from before the damage occurs. This is mainly because correcting for the damage requires an asymmetric configuration in the lattice, which is only possible to a limited degree before interfering with the focus capability of other (undamaged) unit cells. But still, for smaller amounts of damage, reconfiguration allows a return to baseline performance. In a realistic scenario, this could be realized in a "lattice-in-the-loop" setting, meaning that we use the actual lattice to collect where light from each unit cell hits the imaging plane, and then use the measurements in the backward pass [^6] The backward pass is the computational graph used to calculate gradients for the parameters from the cost function. to get the required parameter changes of the physical mirror, i.e., we use the simulation model as a proxy for the error calculation and simply replace predicted values – the distances to the focal point – with real measurements.

### v2

@@ 2.2 Angle-based generalised coordinates
    As the first step, we place the bottom left corner elementary cell of the lattice, which has six degrees of freedom (Fig. 3B, box 1). The six parameters we choose to describe this elementary cell are the location of the lattice and the orientation of the bottom and top beams and levers.
@@
    Then we add the first column, i.e. we stack up elementary cells on top of the first one (Fig. 3B, box 2). Each elementary cell in this column has three degrees of freedom. Here, we choose the orientation of bottom and top lever as well as the orientation of the upper beam as parameters.
@@
    We then add the first row, i.e. we string together elementary cells to the right of the first one (Fig. 3B, box 3). These elementary cells only have two degrees of freedom, which we choose to be the orientation of the bottom beam and lever. In this case, the midpoint of the upper beam ($\mathrm{P^{\prime}}$) is given by the (analytic) solution of a quadratic equation representing the beam/lever constraints (Fig. 3C).
@@
    Then we fill in the remaining elementary cells row-wise from left to right (Fig. 3B, box 4). All these elementary cells have only one degree of freedom, which we choose to be represented by the orientation of the lower lever. The midpoint of the upper beam is again given by the solution of a quadratic equation.
@@
For a detailed mathematical description, see section˜4.2. Although we parametrise the lattice using beam and lever angles here, many other parametrisations are possible. From the quadratic equations, we further derive upper and lower bounds for the lever angles, allowing levers to be pushed or dragged by other lattice elements. For lattices in two dimensions, additional conditions are derived to avoid configurations with overlapping lattice elements. Models of lattices in three dimensions are constructed analogously. For details, see section˜4.3.
Figure 3: Parametrizing Totimorphic lattices. (A) When connecting elementary cells to an existing lattice, overlapping elements are shared with already existing cells (shown in red). Thus, the number of degrees of freedom of the newly added cell is reduced. (B) Iterative construction of a lattice made of Totimorphic cells. (C) If all points except one in an elementary cell are predetermined (shown in blue), the remaining point is obtained from the beam and lever length constraints. This is illustrated here for the case where $\mathrm{A^{\prime}}$ and $\mathrm{C}$ are given. Since $\mathrm{P^{\prime}}$ is connected via a beam and lever to these points, it has to be a distance of $\frac{1}{2}l_{\text{b}}$ away from both. The solution to this problem is given by the intersection points of two circles, or equivalently, by a quadratic equation. Two points $\mathrm{P^{\prime}}$ and $\mathrm{P}^{\prime\prime}$ satisfy the constraints of a Totimorphic lattice, however, the latter would reverse the order of $\mathrm{A^{\prime}}$ and $\mathrm{B^{\prime}}$ and is therefore not a valid physical solution.
2.3 Differentiable model of Totimorphic lattices
[…]
To change the configuration of the lattice to one with some desired property (e.g. a desired stiffness), we minimise an objective-dependent cost function that measures how well the current configuration satisfies the desired property. In most cases, we express the cost function using the physical coordinates obtained from the Totimorphic model. What is actually being optimised, however, are the parameters $\mathcal{P}$ used to characterise the lattice. By tracking intermediate configurations, we obtain the trajectory through parameter space connecting the initial and the final lattice configuration. Each of these configurations is a valid Totimorphic lattice, meaning that the reconfiguration is achieved only through changes in the beam and lever angles (realised via joints), without ever deforming or damaging beams and levers.
In this work, we exclusively use automatic differentiation (i.e. gradient descent) as our optimisation algorithm of choice to guarantee scalability and continuously connected trajectories. In other terms: for real-world prototypes, the control of the actuators (rotation in the joints) is given by the negative gradient of the cost function.
@@ 2.3 Differentiable model of Totimorphic lattices
2.4 First proof of concept: continuous inverse design of lattice structures
@@
As a first proof of concept, we explore reconfigurable lattice structures (on the length scale of cm) with adjustable mechanical properties. We calculate the mechanical properties of a Totimorphic lattice by simulating a compression experiment using the direct stiffness method (see section˜4.4 for details). In a compression experiment, the lattice is glued between surfaces and slowly compressed. During compression, the response of the lattice to this load is measured, for example, the displacement of nodes in the lattice – from which mechanical properties such as effective stiffness and Poisson’s ratio are derived. Here, we restrict the work to the effective Poisson’s ratio, which measures how a material expands on average horizontally when compressed vertically. A positive Poisson’s ratio (like honeycomb tiles) means the material widens, while a negative one means it narrows (also known as auxetic). A Poisson’s ratio of zero consequently means that the material does not change its width during compression (see section˜4.4.3 for details).
2.4.1 Direct stiffness method
@@
For the direct stiffness approach, we only consider beam and lever elements in the lattice, as they are the major supporting elements of the lattice. We model these using Generalised Euler-Bernoulli beam elements [17], described in section˜4.4.1. Furthermore, we assume that joints in the Totimorphic lattice are locked during the compression experiment, meaning that the parameters $\mathcal{P}$ are not changed while applying a load. To do a single step of the compression experiment, we first calculate the stiffness matrix from the lattice points $\mathcal{R}$. Using the stiffness matrix, we can then derive the displacement of each node in the lattice resulting from the compression by solving a linear equation. We repeat this process several times and note down the horizontal and vertical strain (mean length change). Finally, the Poisson’s ratio, which is defined as the slope of the strain-strain curve, is obtained via linear regression. This process can be formalised as a single differentiable function $f_{\mathrm{DS}}$ which takes as input the physical lattice coordinates and returns the Poisson’s ratio $\nu=f_{\mathrm{DS}}\left(f_{\mathrm{T}}(\mathcal{P})\right)$. The trajectory in parameter space ($\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{P}$) to reconfigure the lattice until it has a desired target Poisson’s ratio $\nu^{\mathrm{tgt}}$ is then obtained by descending the gradient of the cost function $C$,
@@
$\frac{\mathrm{d}}{\mathrm{d}t}$ $\mathcal{P}\propto-\nabla_{\mathcal{P}}C(\mathcal{P})\,$, (2)
@@
where $\|\cdot\|_{1}$ is the L1 norm. Further details about the mathematical formulation of the direct stiffness method and the Poisson’s ratio can be found in section˜4.4. The whole optimisation pipeline is illustrated in Fig. 4A.
Figure 4: Inverse design of mechanical properties of Totimorphic lattices. (A) Schematic of the inverse design pipeline. Reconfiguration is driven by gradient descent on a cost function, realized using automatic differentiation (autograd). (B) During optimisation, we obtain a continuous trajectory through configuration space (here represented using the first principal component of the angles representing the lattice) connecting the initial configuration ($\nu=0$) and the two final configurations ($\nu=\pm 0.5$). Since beams and levers are never deformed during reconfiguration, the lattice passes through all intermediate Poisson’s ratios while morphing. (C) Illustration of the final design for $\nu=0.5$ (left) and $\nu=-0.5$ (right), shown without (left) and with load (right, red arrows).
2.4.2 Morphing between auxetic and non-auxetic configurations
@@
We demonstrate this approach for a $4\times 4$ Totimorphic lattice with beam length $l_{\text{b}}=1\,\mathrm{cm}$, beam cross-area $A=4\,\mathrm{mm}^{2}$, Young’s modulus of the beam material of $E=1\,\mathrm{GPa}$ (around the order of magnitude of polymers), and second moment of area $I=\frac{A^{2}}{12}$ (representing a square-shaped cross-sectional area). The initial lattice has a Poisson’s ratio of approximately zero, which is not surprising as it resembles a square lattice. From this initial configuration, we perform gradient descent to get to a configuration with Poisson’s ratio $\nu^{\mathrm{tgt}}=\pm 0.5$. This results in two trajectories starting from the initial configuration and ending at a configuration with the desired Poisson’s ratio, shown in Fig. 4B using the first principal component of the lattice parameters $\mathcal{P}$. The final lattice configurations and their response to applying load is shown in Fig. 4C. In case of the configuration with positive Poisson’s ratio (Fig. 4C, left), large pockets were formed that widen during compression. In case of negative Poisson’s ratio (Fig. 4C, right), the resulting lattice geometry is composed of square-shaped elements (one beam and two levers surrounded by springs) that pivot inwards when compressed from above. This effect is most pronounced in the lower half of the lattice structure. Similar designs were obtained for larger lattices, e.g., $8\times 8$ trained to achieve even higher Poisson’s ratios (Fig. S1 in the Supplemental Information). It is worthy noticing that every intermediate configuration obtained during optimisation is a legitimate Totimorphic lattice, and hence we are capable of continuously adjusting the Poisson’s ratio by continuously changing the lattice angles. For an animation of the reconfiguration process and the compression experiment (after reconfiguration), see Movie S1a,b and Movie S2a,b. Interestingly, the final configurations are not perfectly symmetric. First of all, symmetry is not required to minimise the cost function, so slight asymmetries are not too surprising. However, asymmetries are further caused by (i) the setup of the Totimorphic lattice, where for instance the bottom left node is fixed in space, while the top right node is strongly constrained by the remaining lattice, and (ii) limited numerical precision in the forward calculation of lattice points. Further simulation details are in section˜4.6.
@@
2.5 Second proof of concept: deployable and reconfigurable large-scale structures
As a second proof of concept, we present a telescope mirror with reconfigurable surface shape, evaluated using ray-based optics simulations. We envision that the Totimorphic lattice (on the length scale of meters) acts as a skeleton for, e.g., a graphene sheet that forms the reflective surface of a primary mirror of a space telescope [20], or as a reconfigurable support truss for antennas [26, 30]. We assume a $6\times 6$ Totimorphic lattice with $l_{\text{b}}=1$m in its flat surface configuration.
2.5.1 Mirror deployment
@@
By only changing the lever angle of all elements in the lattice (i.e., setting them all to the same value), we can switch from the flat surface configuration to a collapsed one (and vice versa), greatly reducing the surface area and volume required to launch the Totimorphic structure (see Fig. 5). The reduction in surface area is given by $\sin\theta_{\mathrm{min}}$, where $\theta_{\mathrm{min}}$ is the minimum angle the levers can be set to physically. Consequently, for launching, the lattice is configured in its collapsed configuration and for deployment, the lever angles are slowly increased to unfold the lattice into a flat surface.
2.5.2 Continuously adjusting the mirror’s focal point
After deployment, the lattice configuration is continuously adjustable to change the optical properties of its mirror surface. In our case, the goal is to focus the light onto a secondary optical element such as a detector or a secondary mirror located at point $\boldsymbol{f}^{\mathrm{tgt}}$, which we call the target focal point in the following (Fig. 6A). We utilise a similar approach as for the material inverse design use-case, meaning that we construct a differentiable framework that allows us to find suitable lattice configurations (and trajectories connecting them) using gradient-based optimisation. First, for simplicity, we assume that each unit cell reflects light at the mid-point of the lever in simulations (i.e., the mid-point between $\mathrm{C}$ and $\mathrm{P}$ as well $\mathrm{C}$ and $\mathrm{P^{\prime}}$, see Fig. 6A). In addition, we assume that the object to be imaged is far away such that the incoming light is parallel. Given an incidence direction $\boldsymbol{I}$ ($\|\boldsymbol{I}\|_{2}=1$, where $\|\cdot\|_{2}$ is the Euclidean norm) of the incoming light, the direction of the reflected light originating from each unit cell $(ij)$ is calculated using ray-based optics. Subsequently, we obtain the point where each reflected light ray is closest to the target focal point (see section˜4.5 for details). The whole calculation can again be formulated as a series of differentiable operations, which we summarise in two functions: the Totimorphic model ($f_{\mathrm{T}}$), and the optics part $f_{\mathrm{O}}$, such that $\boldsymbol{F}=f_{\mathrm{O}}(f_{\mathrm{T}}(\mathcal{P}))$ is a list of vectors containing the closest point to $\boldsymbol{f}^{\mathrm{tgt}}$ on each reflected light ray.
@@ 2.5.2 Continuously adjusting the mirror’s focal point
Figure 6: Deployment and reconfiguration of a Totimorphic telescope mirror. (A) Schematic of the studied setup. Given an object to be imaged, the unit cells of the Totimorphic lattice are used as mirror elements to focus the light on an imaging plane. By changing the lattice configuration, the focal point and focal length are adjusted.(B) Reconfiguration of the lattice to several different focal lengths, starting from a flat configuration. A few parameters as well as the objective-dependent cost are shown. (C) Lattice configuration and images received when using the lattice as a mirror, for configurations featuring different focal lengths. Credit for image of Pluto: NASA/JHUAPL/SwRI.
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{P}\propto$ $-\nabla_{\mathcal{P}}C(\mathcal{P})\,$, (4)
During the optimisation process, we only present light coming from one direction (i.e., orthogonal to the initial lattice surface) which has to be focused into a single target point. In Fig. 6B-C, we show how the Totimorphic lattice transforms – starting from a flat surface – into curved surfaces that focus the light into a single point at decreasing focal lengths given in multiples of the mirror diameter $D$. Fig. 6B (bottom) shows the value of the cost function during deployment (from flat to a focal length of 20DD, see Fig. S2 in the Supplemental Information for further illustration) and during reconfiguration to new targets, i.e., focal lengths of 16, 12, 8, 4, 2, and 1DD. In all cases, we reach the same cost after reconfiguring the lattice, which means that the light reflected from all unit cells is focused with a similar precision in the target focal points. In Fig. 6B (top), a few lattice parameters are shown during the reconfiguration process, demonstrating that only minor continuous changes are required to achieve the desired result. The reconfiguration is further illustrated in Movie S3a,b.
As a proof of concept, we use the Totimorphic lattice to image a real object; specifically, a photograph of Pluto at the same distance and angular extension of the Moon (as seen from Earth), see Fig. 6A. To construct the image, we record where light coming from each pixel of the image hits the imaging plane after being reflected by the unit cells of the lattice – imitating the real process of image generation in a telescope (see section˜4.5.4 for details). The resulting image closely resembles the original object, resolving small details such as contours of icy regions on Pluto (Fig. 6C). Moreover, as expected, decreasing the focal length of the Totimorphic mirror increases its field of view; enabling adaptive zoom-in and out by reconfiguring the lattice.
Although we only demonstrate reconfiguration for a few discrete focal lengths, in principle, the telescope can achieve a continuum of focal points by successively adjusting the target focal length. In addition, the telescope is not limited to focal points that are on the central axis of the telescope surface, as shown in Fig. 7A-B where we guide the focal point on a circle around this axis using the above gradient-based optimisation.
@@
2.5.3 Self-repair
The presented approach also allows to use reconfiguration of the lattice surface to compensate for defects introduced by mirror damage, e.g., through micrometeoroids [21]. We illustrate this by adding a single defect to one of the lattice’s unit cells (selected randomly), which is modelled by adding a deflection (randomly determined but static) to the light rays reflected from this cell (Fig. 7C; see section˜4.5.3 for details). As before, we perform gradient descent on the respective cost function to find a lattice configuration that compensates for this defect (Fig. 7C-D). This is studied for increasing degrees of damage, i.e., for increasing amounts of deflection. For all cases, the reconfiguration improves the performance of the mirror – although for more extreme damage, we are not able to return to the baseline performance from before the damage occurs. This is mainly because correcting for the damage requires an asymmetric configuration in the lattice, which is only possible to a limited degree before interfering with the focus capability of other (undamaged) unit cells. But still, for smaller amounts of damage, reconfiguration allows a return to baseline performance. In a realistic scenario, this could be realised in a “lattice-in-the-loop” setting, meaning that we use the actual lattice to collect where light from each unit cell hits the imaging plane, and then use the measurements in the backward pass (i.e., the computational graph used to calculate gradients for the parameters from the cost function) to get the required parameter changes of the physical mirror. Or put simply: using the simulation model as a proxy for the error calculation and replacing predicted values – the distances to the focal point – with real measurements. The same approach can be applied for other damage types as well, such as damaged actuators which would be modelled by keeping the corresponding generalised coordinate fixed or by adding strong penalty terms to the cost function. Further simulation details are in section˜4.7.